UT-Thesis Class File Example

by

François Pitt

A thesis submitted in conformity with the requirements

for the degree of Doctor of Philosophy Graduate Department of Computer Science

University of Toronto

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1999 by François Pitt

Abstract

UT-Thesis Class File Example

François Pitt Doctor of Philosophy

Graduate Department of Computer Science

University of Toronto 1999

Contents

Bibliography 1

Bibliography

1